# Addressing divergent representations from causal interventions on neural networks

### v3

@@ Abstract
A common approach to mechanistic interpretability is to causally manipulate model representations via targeted interventions in order to understand what those representations encode. Here we ask whether such interventions create out-of-distribution (divergent) representations, and whether this raises concerns about how faithful their resulting explanations are to the target model in its natural state. First, we demonstrate theoretically and empirically that common causal intervention techniques often do shift internal representations away from the natural distribution of the target model. Then, we provide a theoretical analysis of two classes of such divergences: ‘harmless’ divergences that occur in the null-space of the weights and from covariance within behavioral decision boundaries, and ‘pernicious’ divergences that activate hidden network pathways and cause dormant behavioral changes. Finally, in an effort to mitigate the pernicious cases, we apply and modify the Counterfactual Latent (CL) loss from Grant (2025) allowing representations from causal interventions to remain closer to the natural distribution, reducing the likelihood of harmful divergences while preserving the interpretive power of the interventions. Together, these results highlight a path towards more reliable interpretability methods.
 [^1] Link to public repository: https://github.com/grantsrb/rep_divergence
@@ 1 Introduction
In this work, we provide both empirical and theoretical insight on these issues. We first demonstrate that divergent representations are a common issue for causal interventions – across a wide range of intervention methods, we find that intervened representations often do diverge from the target NN’s natural distribution. We then provide theoretical examples of two types of divergence: ‘harmless’ divergences that can occur from within-decision-boundary covariance along causal dimensions or from deviations in the null-spaces of the NN layers, and ‘pernicious’ divergences that activate hidden network pathways and can cause dormant changes to behavior. We provide discussion on how harmless and pernicious cases are not always mutually exclusive, where the harm often depends on the specific mechanistic claims. Finally, we provide a broad-stroke, initial solution for mitigating pernicious divergences by minimizing all intervened divergences. We show that we can use the Counterfactual Latent (CL) auxiliary loss introduced in grant2025mas to reduce all representational divergence in the Boundless DAS setting from wu2024alpacadas while maintaining the same behavioral accuracy; and we introduce a modified version of the CL loss that targets causal subspaces and show that it can improve out-of-distribution (OOD) intervention performance on synthetic tasks. Although we do not propose this method as the final solution to representational divergence, we pose it as a step towards more reliable interventions.
@@
    We provide a theoretical treatment of cases in which divergence can arise innocuously from variation in the null-space and from covariance within decision boundaries of behavioral subspaces, demonstrating that some divergences can be okay and even desired (Section 4.1)
@@
    We use the CL auxiliary loss from grant2025mas to minimize patching divergence directly in the Large Language Model (LLM) Boundless DAS experimental setting from wu2024alpacadas, and we introduce a stand-alone, modified CL loss that exclusively minimizes divergence along causal dimensions, improving OOD intervention performance in synthetic settings. Together, these results provide an initial step towards preventing pernicious divergences (Section 5).
@@ 4 When are divergent representations harmless or pernicious?
Having demonstrated that divergent representations are a common phenomenon, we now turn our attention to considering whether divergence is a concerning phenomenon. We propose that divergence is harmless to most functional claims if it occurs in the behavioral null-space of the model weights or from covariance within decision boundaries, and that it can be pernicious if it recruits hidden pathways or leads to dormant behavioral changes. However, we stress that the harm is inherently claim dependent, meaning that these divergences are not mutually exclusive.
@@ 4.1 Harmless cases
We first explore cases in which we might consider divergent representations to be harmless to many general functional claims. First among these is a set of cases in which divergence is bottlenecked into the null-space of the next computational layer(s). Formally, we define the null-space of a weight matrix $W\in\mathbb{R}^{d_{2}\times d_{1}}$ as $\mathcal{N}(W)=\{v\in\mathbb{R}^{d_{1}}\,\,|\,\,Wv=0\}$. Neural activity in the null-space of the weights refers to any changes $\delta\in\mathbb{R}^{d_{1}}$ for which $W(h^{\ell}+\delta)=Wh^{\ell}$. We can generalize this notion beyond the weights to all remaining computations after and including layer $\ell$ in a neural network denoted $f_{\geq\ell}(h)$. We call this the behavioral null-space for a layer $\ell$, defined as $\mathcal{N}(f_{\geq\ell}(h^{\ell}))=\{v\in\mathbb{R}^{d_{1}}\,\,|\,\,f_{\geq\ell}(h^{\ell}+v)=f_{\geq\ell}(h^{\ell})\}$.
If one cares only about the model’s output behavior, then divergence that is bottlenecked into null-spaces and that therefore has no effect on the model’s behavior would be harmless. Indeed, this covers many present day scenarios, and one could even desire to intervene along null-space directions in order to causally test that such dimensions are indeed in the null-space. It is, however, difficult to define the full behavioral null-space in practice due to the possibility of dormant behavioral changes, which we discuss in Section 4.2.2.
A less obvious case of harmless divergence can occur from covariance within a network’s decision boundaries. For example, the top panel of Figure 1(b) provides a simulated depiction of harmless divergence caused by covariance along causal dimensions. Here, divergent representations are created by swapping the first causal dimension of sampled natural vectors, however, these representations remain within the desired decision boundaries, resulting in the appropriate counterfactual behavior.
We can expand this intuition using a behaviorally binary subspace as an example, where we define a behaviorally binary subspace as a vector subspace which causally impacts the outputs of a future processing layer (e.g., classification labels) only through its sign. Formally, let $f:\mathbb{R}^{d}\to\mathbb{R}^{k}$ denote a computational layer (possibly consisting of multiple NN layers and functions) where $k$ is the output dimensionality. Let $\operatorname{sign}(\cdot)$ denote the elementwise sign map $\operatorname{sign}:\mathbb{R}^{d_{\text{var}}}\to\{-1,1\}^{d_{\text{var}}}$, and assume a fixed alignment function $\mathcal{A}$ and subspace selection matrix $D_{\text{var}}\in\mathbb{R}^{d\times d}$. A linear subspace $Z\subseteq\mathbb{R}^{d}$ is behaviorally binary (with respect to $f$, $D_{\text{var}}$, and $\mathcal{A}$) iff for all $D_{\text{var}}\mathcal{A}(h),D_{\text{var}}\mathcal{A}(h^{\prime})\in Z$,
@@
Now, suppose we have an NN with two causal subspaces, $\mathbf{\vec{z}_{\text{var}_{a}}}\subseteq\mathbb{R}^{d_{\text{var}_{a}}}$ and $\mathbf{\vec{z}_{\text{var}_{b}}}\subseteq\mathbb{R}^{d_{\text{var}_{b}}}$, with values $\vec{z}^{\,(x_{i})}_{\text{var}_{a}}$} and $\vec{z}^{\,(x_{i})}_{\text{var}_{b}}$}, for a model input $x_{i}$, where we use the bold, tilde notation to distinguish variables from their values. Furthermore, assume that $\mathbf{\vec{z}_{\text{var}_{b}}}$ is a behaviorally binary subspace that co-varies with, $\mathbf{\vec{z}_{\text{var}_{a}}}$. Using $h^{(x_{i})}$ and $z^{(x_{i})}$ from Equation 1 under a given input $x_{i}$, we use the following definition:
@@
Due to the assumption of covariance in $\mathbf{\vec{z}_{\text{var}_{a}}}$ and $\mathbf{\vec{z}_{\text{var}_{b}}}$, it is reasonable to assume that the values $\vec{z}^{\,(x_{\text{low}})}_{\text{var}_{b}}$} and $\vec{z}^{\,(x_{\text{high}})}_{\text{var}_{b}}$} are systematically distinct for distinct values of $\mathbf{\vec{z}_{\text{var}_{a}}}$ under some classes of inputs, $x_{\text{low}}$ and $x_{\text{high}}$, while $\operatorname{sign}({\vec{z}^{\,(x_{\text{low}})}_{\text{var}_{b}}})=\operatorname{sign}({\vec{z}^{\,(x_{\text{high}})}_{\text{var}_{b}}})$. Under these assumptions, if we perform an interchange intervention on $\mathbf{\vec{z}_{\text{var}_{b}}}$ using source representations from input $x_{\text{low}}$ and target representations from input $x_{\text{high}}$, the intervened representation will have values:
@@
Because we assumed that the value of $\vec{z}^{\,(x_{\text{low}})}_{\text{var}_{b}}$ is systematically unique due to covariance in $\mathbf{\vec{z}_{\text{var}_{a}}}$ and $\mathbf{\vec{z}_{\text{var}_{b}}}$, the values $\vec{z}^{\,(x_{\text{high}})}_{\text{var}_{a}}$ and $\vec{z}^{\,(x_{\text{low}})}_{\text{var}_{b}}$ in Equation 6 will have never existed together in the native distribution, but the behavior of the NN will remain unchanged because $\mathbf{\vec{z}_{\text{var}_{b}}}$ is behaviorally binary and its sign has not changed. This is thus a case of harmless divergence.
If we focus only on the divergence in causal subspaces and ignore cases of null-space interactions, what claims can we make about the model’s natural neural mechanisms using divergent, intervened representations? An interpretation consistent with the principles of superposition (Smolensky1988; PDP1; PDP2; elhage2022superposition) is that differences in the exact values of each causal subspace do not matter – rather, only the decision boundaries along these subspaces matter. Under this interpretation, any divergence in the intervened distribution arising from covariance in causal subspaces is acceptable because the separation of the causal subspaces and their decision boundaries (i.e. the computationally important aspects of the NN) are respected in the alignment. This interpretation allows us to ignore computationally irrelevant covariance in our attempt to understand NN mechanisms.
@@ 4.2 Pernicious divergence via off-manifold activation
Pernicious cases of representational divergence involve the concept of hidden pathways, which refer to any unit, vector direction, or subcircuit that is inactive on the natural support of representations for a given context but becomes active and influences behavior under an intervention. Although hidden pathways can be compatible with harmless divergences (Section 4.1), they can also undermine claims about natural mechanisms and can prime dormant behavioral changes discussed in Section 4.2.2.
Formally, let $\mathcal{D}$ denote the data distribution over latent representations $h^{\ell}\in\mathbb{R}^{d}$ at layer $\ell$, and let $\mathcal{S}=\mathrm{supp}(h^{\ell}\sim\mathcal{D})$ denote its support. Denoting the intended class $K$ following an intervention as subscripted $\rightarrow K$, an intervened representation $\hat{h}^{\ell}_{\rightarrow K}$ is said to be divergent if $\hat{h}^{\ell}_{\rightarrow K}\notin\mathcal{S}$ or if it causes a ReLU state change for any neuron that does not change sign for the same class K under $\mathcal{S}$. We define the convex hull of class-K representations as $\mathrm{conv}(S_{K})=\{\sum_{i}\alpha_{i}h_{i,K}^{\ell}:\alpha_{i}\geq 0,\,\sum_{i}\alpha_{i}=1\}$ where the subscript $K$ denotes that $h^{\ell}_{i,K}$ was taken from class $K$ inputs. Projecting an intervention onto $\mathrm{conv}(S_{K})$ ensures that it remains within the convex interpolation region of class-K samples.
@@ 4.2.1 Mean-difference patching can activate hidden pathways
so $s_{A}\in\{0.25,0.5\}$. For class B, all outputs are zero, $s_{B}=0$.
@@
Applying this to class B representations gives
@@
The intervention flips the decision to class A ($\hat{s}>0$). However, the third hidden unit becomes active only for $\hat{h}^{\ell}_{B\rightarrow A}$, never for natural $h_{A}^{\ell}$. This new activation is a hidden pathway that was silent under all native samples. Thus the mean-difference patch crosses the decision boundary only by activating an off-manifold circuit.
If we project $\hat{h}^{\ell}_{B\rightarrow A}$ onto $\mathrm{conv}(S_{A})$ (or equivalently onto a local PCA subspace of $S_{A}$), this ReLU state change disappears, and the decision boundary is no longer crossed—confirming that the original effect was driven by an off-manifold activation rather than a genuine within-manifold causal mechanism. This experiment does not support claims about the natural mechanisms of the model.
@@ 4.2.2 Dormant behavioral changes
Divergent representations can also yield dormant behavioral changes—interventions that appear behaviorally neutral in one context but alter predictions in another. This occurs when an off-manifold activation primes a downstream computation that only manifests under specific contextual inputs.
@@
Extend the network from the previous Section 4.2.1 by adding one row of zeros to $W_{\ell}$ and $b_{\ell}$, producing $h^{\ell+1}\in\mathbb{R}^{4}$ where the final coordinate is always zero. Add a context vector $v\in\mathbb{R}^{4}$ and a final affine layer:
@@
The model predicts class A when $v_{4}<0.75$ but switches to class C when $v_{4}>0.75$. Notably, $v_{4}$ would not naturally cause a class C prediction below a value of $1$ due to the bias threshold. Thus the same intervention that was benign in one context ($v_{4}<0.75$) produces a behavioral flip in another ($1>v_{4}>0.75$) purely due to the latent divergence priming a new pathway.
@@
Hidden pathways arise when off-manifold representations activate computations that never occur for natural representations. Such pathways can alter causal conclusions even when behavior appears unchanged depending on the claims. Manifold-preserving projections and ReLU-pattern audits provide practical safeguards against these pernicious forms of divergence.

### v4

@@ Abstract
A common approach to mechanistic interpretability is to causally manipulate model representations via targeted interventions in order to understand what those representations encode. Here we ask whether such interventions create out-of-distribution (divergent) representations, and whether this raises concerns about how faithful their resulting explanations are to the target model in its natural state. First, we demonstrate theoretically and empirically that common causal intervention techniques often do shift internal representations away from the natural distribution of the target model. Then, we provide a theoretical analysis of two cases of such divergences: ‘harmless’ divergences that occur in the behavioral null-space of the layer(s) of interest, and ‘pernicious’ divergences that activate hidden network pathways and cause dormant behavioral changes. Finally, in an effort to mitigate the pernicious cases, we apply and modify the Counterfactual Latent (CL) loss from Grant (2025) allowing representations from causal interventions to remain closer to the natural distribution, reducing the likelihood of harmful divergences while preserving the interpretive power of the interventions. Together, these results highlight a path towards more reliable interpretability methods.  [^1] Link to public repository: https://github.com/grantsrb/rep_divergence
@@ 1 Introduction
In this work, we provide both empirical and theoretical insight on these issues. We first demonstrate that divergent representations are a common issue for causal interventions – across a wide range of intervention methods, we find that intervened representations often do diverge from the target NN’s natural distribution. We then provide theoretical examples of two types of divergence: ‘harmless’ divergences that can occur from within-decision-boundary covariance along causal dimensions or from deviations in the null-spaces of the NN layers, and ‘pernicious’ divergences that activate hidden network pathways and can cause dormant changes to behavior. We provide discussion on how harmless and pernicious cases are not always mutually exclusive, where the harm depends on the specific mechanistic claims. Finally, we provide a broad-stroke, initial solution for mitigating pernicious divergences by minimizing all intervened divergences. We show that we can use the Counterfactual Latent (CL) auxiliary loss introduced in grant2025mas to reduce all representational divergence in the Boundless DAS setting from wu2024alpacadas while maintaining the same behavioral accuracy; and we introduce a modified version of the CL loss that targets causal subspaces and show that it can improve out-of-distribution (OOD) intervention performance on synthetic tasks. Although we do not propose this method as the final solution to representational divergence, we pose it as a step towards more reliable interventions.
@@
    We provide a theoretical treatment of cases in which divergence can arise innocuously from variation in null-spaces, demonstrating that some divergences can be okay and even desired (Section 4.1)
@@
    Lastly, we use the CL auxiliary loss from grant2025mas to minimize patching divergence directly in the 7B Large Language Model (LLM) Boundless DAS experimental setting from wu2024alpacadas, and we introduce a stand-alone, modified CL loss that exclusively minimizes divergence along causal dimensions, improving OOD intervention performance in synthetic settings. Together, these results provide an initial step towards mitigating pernicious divergences (Section 5).
@@ 4 When are divergent representations harmless or pernicious?
Having demonstrated that divergent representations are a common phenomenon, we now consider whether divergence is a concerning phenomenon. We propose that divergence is harmless to many functional claims if it exists in the behavioral null-space, and that it can be pernicious if it recruits hidden pathways or causes dormant behavioral changes. However, we stress that the harm is inherently claim dependent, meaning that these forms of divergence are not mutually exclusive.
@@ 4.1 Harmless cases
This section explores a set of cases for which we might consider divergent representations to be harmless to functional claims. We first focus on cases in which divergence is bottle-necked into the null-spaces of the next interacting weight matrices. Formally, we define the null-space of a weight matrix $W\in\mathbb{R}^{d^{\prime}\times d}$ as $\mathcal{N}(W)=\{v\in\mathbb{R}^{d}\,\,|\,\,Wv=0\}$. Neural activity in the null-space of the weights refers to any changes $\delta\in\mathbb{R}^{d}$ for which $W(h+\delta)=Wh$. We propose that divergence $v\in\mathcal{N}(W)$ is harmless to the computation of $W$ because it is equivalent to adding the zero vector $W(h+v)=Wh+Wv=Wh+0=W(h+0)$. Notably, however, this harmlessness does not apply to the sub-computations of the matrix multiplication because $v\in\mathcal{N}(W)$ does not imply that $W_{i,j}(h_{j}+v_{j})=W_{i,j}(h_{j}+0)$ for vector row $j$ and matrix row $i$. Thus, $v$ in this case is potentially harmful to mechanistic claims about individual activation-weight sub-computations, while being harmless to the overall matrix multiplication.
We can generalize this notion beyond matrices to an arbitrary function $\psi$. Let $\psi:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$ and let $X\subseteq\mathbb{R}^{d}$. We define the behavioral null-space with respect to $X$ as
$\mathcal{N}(\psi,X)=\{\,v\in\mathbb{R}^{d}\mid\forall x\in X,\;\psi(x+v)=\psi(x)\,\}.$ (4)
A common case of $\psi$ in practice for a layer $\ell$ of an NN $f$ consists of all subsequent computations after and including layer $\ell$, denoted $f_{\geq\ell}(h)$. We propose that behaviorally null divergence $v$ is harmless to the overall computation of $f_{\geq\ell}$ because it is equivalent to adding 0 to the input. However, $v$ can be harmful to claims about a sublayer $\ell+k$ within $f_{\geq\ell}$ because $f_{\ell+k}(f_{\geq\ell,<\ell+k}(h+v))$ is not guaranteed to be equal to $f_{\ell+k}(f_{\geq\ell,<\ell+k}(h+0))$ (Sec. 4.2.1). See Appx. A.3 and Algorithm 1 to practically classify harmlessness when $\mathcal{N}(\psi,X)$ characterizes the full space of harmless divergence.
Idealized Case Study. We now present an example of harmless divergence in the behavioral null-space by introducing a behaviorally binary subspace—a vector subspace which causally impacts the outputs of a future processing layer (e.g., classification labels) only through its sign. Formally, let $f:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$ denote a computational unit (possibly consisting of multiple NN layers and functions). Let $\operatorname{sign}(\cdot)$ denote the elementwise sign map $\operatorname{sign}:\mathbb{R}^{d_{\text{var}}}\to\{-1,1\}^{d_{\text{var}}}$, and assume a fixed alignment function $\mathcal{A}$ and subspace selection matrix $D_{\text{var}}\in\mathbb{R}^{d\times d}$ (Sec. 2.2). A linear subspace $Z\subseteq\mathbb{R}^{d}$ is behaviorally binary (with respect to $f$, $D_{\text{var}}$, and $\mathcal{A}$) iff for all $D_{\text{var}}\mathcal{A}(h),D_{\text{var}}\mathcal{A}(h^{\prime})\in Z$,
@@
Now, suppose we have an NN with two causal subspaces, $\mathbf{\vec{z}_{\text{var}_{a}}}\subseteq\mathbb{R}^{d_{\text{var}_{a}}}$ and $\mathbf{\vec{z}_{\text{var}_{b}}}\subseteq\mathbb{R}^{d_{\text{var}_{b}}}$, with values $\vec{z}^{\,(x_{i})}_{\text{var}_{a}}$} and $\vec{z}^{\,(x_{i})}_{\text{var}_{b}}$}, for a model input $x_{i}$, where the bold, tilde notation distinguishes variables from their values. Further assume that $\mathbf{\vec{z}_{\text{var}_{b}}}$ is behaviorally binary and co-varies with, $\mathbf{\vec{z}_{\text{var}_{a}}}$. Using $h^{(x_{i})}$ and $z^{(x_{i})}$ from Equation 1 under a given input $x_{i}$, we use the following definition:
@@
Due to the assumption of covariance in $\mathbf{\vec{z}_{\text{var}_{a}}}$ and $\mathbf{\vec{z}_{\text{var}_{b}}}$, it is reasonable to assume that the values $\vec{z}^{\,(x_{\text{low}})}_{\text{var}_{b}}$} and $\vec{z}^{\,(x_{\text{high}})}_{\text{var}_{b}}$} are systematically distinct for distinct values of $\mathbf{\vec{z}_{\text{var}_{a}}}$ under some classes of inputs, $x_{\text{low}}$ and $x_{\text{high}}$, while $\operatorname{sign}({\vec{z}^{\,(x_{\text{low}})}_{\text{var}_{b}}})=\operatorname{sign}({\vec{z}^{\,(x_{\text{high}})}_{\text{var}_{b}}})$. Now, an interchange intervention on $\mathbf{\vec{z}_{\text{var}_{b}}}$ using a source vector from $x_{\text{low}}$ and target vector from $x_{\text{high}}$, will produce:
@@
Because we assumed that the value of $\vec{z}^{\,(x_{\text{low}})}_{\text{var}_{b}}$ is systematically unique due to covariance in $\mathbf{\vec{z}_{\text{var}_{a}}}$ and $\mathbf{\vec{z}_{\text{var}_{b}}}$, the values $\vec{z}^{\,(x_{\text{high}})}_{\text{var}_{a}}$ and $\vec{z}^{\,(x_{\text{low}})}_{\text{var}_{b}}$ in Equation 7 will have never existed together in the native distribution, but the behavior will remain unchanged because $\mathbf{\vec{z}_{\text{var}_{b}}}$ is behaviorally binary and its sign has not changed. This divergence is thus harmless to the claim of discovering $f$’s causal axes.
Summary. We propose that divergences within behavioral null-spaces are harmless to many functional claims about a function $\psi$’s computations when the claim encapsulates (i.e. ignores) the internal, sub-computations of $\psi$. However, we are not suggesting that the behavioral null-space encompasses the set of all harmless divergences. In general, such an exhaustive set cannot exist without assuming the superiority of some scientific claims/assumptions over others. For example, take the set of all harmless divergences for a specific claim, then modify the claim to assume it permissible to diverge in a manner previously excluded from the harmless set. The modification to the claim also modifies the harmless divergences. Lastly, we note that behaviorally null divergence is not always harmless. Indeed, one could even desire to intervene on the behavioral null-space to causally test that it is null. Thus, we stress that the mechanistic claim for which an intervention is meant to support is important for determining the harmlessness of the divergence.
@@ 4.2 Pernicious divergence via off-manifold activation
We now explore pernicious cases of representational divergence involving the concept of hidden pathways, which refer to any unit, vector direction, or subcircuit that is inactive on the natural support of representations for a given context but becomes active and influences behavior under an intervention. Although hidden pathways can be compatible with harmless divergences (Sec. 4.1), they can also undermine claims about natural mechanisms and can prime dormant behavioral changes discussed in Sec. 4.2.2.
Formally, let $\mathcal{D}$ denote the data distribution over latent representations $h^{\ell}\in\mathbb{R}^{d}$ at layer $\ell$, and let $\mathcal{S}=\mathrm{supp}(h^{\ell}\sim\mathcal{D})$ denote its support with $\mathcal{S}_{K}=\mathrm{supp}(h^{\ell}_{K}\sim\mathcal{D})$ the support for class $K$. Denoting the intended class $K$ following an intervention as subscripted $\rightarrow K$, an intervened representation $\hat{h}^{\ell}_{\rightarrow K}$ is said to be divergent if $\hat{h}^{\ell}_{\rightarrow K}\notin\mathcal{S}_{K}$ (e.g. it exists off the natural manifold of class $K$). We define the convex hull of class-K representations as $\mathrm{conv}(S_{K})=\{\sum_{i}\alpha_{i}h_{i,K}^{\ell}:\alpha_{i}\geq 0,\,\sum_{i}\alpha_{i}=1\}$ where the subscript $K$ denotes that $h^{\ell}_{i,K}$ was taken from class $K$ inputs. Projecting an intervention onto $\mathrm{conv}(S_{K})$ ensures that it remains within the convex interpolation region of class-K.
@@ 4.2.1 Mean-difference patching can activate hidden pathways
so for class-A, $s_{A}\in\{0.25,0.5\}$. For class-B, all outputs are zero, $s_{B}=0$.
@@
Applying this to class-B representations gives
@@
The intervention flips the decision to class-A ($\hat{s}>0$). However, the third hidden unit becomes active only for $\hat{h}^{\ell}_{B\rightarrow A}$, never for natural $h_{A}^{\ell}$. This new activation is a hidden pathway that was silent under all native samples. Thus the mean-difference patch crosses the decision boundary only by activating an off-manifold circuit.
If we project $\hat{h}^{\ell}_{B\rightarrow A}$ onto $\mathrm{conv}(S_{A})$ (or equivalently onto a local PCA subspace of $S_{A}$), this ReLU state change disappears, and the decision boundary is no longer crossed—confirming that the original effect was driven by divergence rather than a within-manifold causal mechanism. It is unclear what this patching experiment reveals about the natural mechanisms of the model.
@@ 4.2.2 Dormant behavioral changes
Divergent representations can also yield dormant behavioral changes: perturbations that appear behaviorally null in one subset of contexts but alter behavior in others. Formally, let $\psi:\mathbb{R}^{d}\times\mathcal{C}\to\mathbb{R}^{d^{\prime}}$ and let $\mathcal{C}_{1}\subset\mathcal{C}$ be a subset of contexts. The space of dormant behavioral changes relative to $X,\mathcal{C}_{1},\mathcal{C}$ is $\mathcal{V}(\psi,X,\mathcal{C}_{1},\mathcal{C})=\mathcal{N}(\psi,X,\mathcal{C}_{1})\setminus\mathcal{N}(\psi,X,\mathcal{C}).$
@@
Extend the network from the previous Sec. 4.2.1 by adding one row of zeros to $W_{\ell}$ and $b_{\ell}$, producing $h^{\ell+1}\in\mathbb{R}^{4}$ where the final coordinate is always zero. Add a context vector $v\in\mathbb{R}^{4}$ and a final affine layer:
@@
The model predicts class-A when $v_{4}<0.75$ but switches to class-C when $v_{4}>0.75$. Notably, $v_{4}$ would not naturally cause a class-C prediction below a value of $1$ due to the bias threshold. Thus the same intervention that was benign in one context ($v_{4}<0.75$) produces a behavioral flip in another ($1>v_{4}>0.75$) purely due to the latent divergence priming a new pathway.
@@
Hidden pathways arise when off-manifold representations activate computations that never occur for natural representations. Such pathways can potentially alter causal conclusions even when behavior appears unchanged. Manifold-preserving projections and ReLU-pattern audits are potentially practical safeguards against these pernicious forms of divergence.